OpenStudy (anonymous):

Find the area and the circumference of a circle with radius of 5. Use the value 3.14 for pie and round your answers.

OpenStudy (raden):

use the formula : area = pi r^2 circumference = 2pi r

OpenStudy (anonymous):

am i multiplying the radius to get my answer? And am i multiplying 2 times 3.14 times 5 to get my answer for the circumference@RadEn

OpenStudy (raden):

for to get its area, we multiplying the radius with r squared, it means : area = 3.14 * 5^2 = 3.14 * 5 * 5 = ....

OpenStudy (anonymous):

i got 78.5 for the radius @RadEn

OpenStudy (raden):

yes, that right... that is the area of circle

OpenStudy (anonymous):

so if i round it would be 78 right @RadEn

OpenStudy (raden):

yeah... btw, what about u with circumference, did u got it ?

OpenStudy (raden):

ur steps were right, multiplying 2 times 3.14 times 5 C = 2 * 3.14 * 5 = ...

OpenStudy (anonymous):

okay but i need area

OpenStudy (anonymous):

@RadEn

OpenStudy (raden):

u got its area = 78.5, right ? :)

OpenStudy (anonymous):

but what is the circumference ? @RadEn

OpenStudy (raden):

C = 2 * 3.14 * 5 = ... ?

OpenStudy (anonymous):

31.4 i got @RadEn

OpenStudy (raden):

yup, that's right

OpenStudy (anonymous):

is it m2 or m3 units? @RadEn

OpenStudy (raden):

actually, for area its unit use m2 and circumference just m not m2 or m3